# Saumili Janaa,1, John Kolinskib,2, Detlef Lohsea,c,3, and Vatsal Sanjaye,4

### v2

@@ 4.3 Predictive Interpolating Model for Maximum Impact Force
For intermediate conditions ($El\sim\mathcal{O}\left(1\right)$), the sphere’s deformation and the fluid’s inertia both contribute, and the peak force deviates from either pure Hertz or Wagner scaling alone. We therefore express the dimensionless peak force as a weighted transition between the two asymptotic contributions following the approach of Sanjay & Lohse (2025)sanjayUnifyingTheoryScaling2025,
@@
with a smooth transition function $\text{f}\left(El\right)$ defined as a tanh function in $\log_{10}El$, centered at $El\sim\mathcal{O}\left(1\right)$. The coefficients are fixed a priori: the Hertz prefactor 5.3 follows directly from the elastic analysis above (without fitting), while the constants 3.24 (Wagner plateau) and 3.2 (the $We^{-1}$ capillary correction) are taken from the Newtonian impact model of Sanjay & Lohse (2025)sanjayUnifyingTheoryScaling2025. This construction ensures a continuous interpolation between the Hertz and Wagner limits. In spirit it follows the additive scaling approach of Sanjay & Lohse (2025)sanjayUnifyingTheoryScaling2025 for drop impacts, but unlike their model – which includes separate viscous regimes – here only the two primary regimes (elastic vs. inertial) are needed. The resulting formula smoothly bridges the two asymptotes and correctly reproduces the peak-force scaling in both limits (this predictive curve is plotted in fig. 4a for comparison).
@@ 5 Influence of Elastic Stress Relaxation
The results in § 3 established the two asymptotic force scalings for impacts with permanent memory ($De\to\infty$): a Wagner plateau at small elasticity number and a Hertz law at large elasticity number (figure 4). These two limits also bound impacts when the material memory is finite: relaxing the memory shifts the response continuously from solid‑like to liquid‑like, with the transition controlled by the non‑dimensional relaxation time. We quantify the memory effect with the Weissenberg number, $Wi$ (eq. 8), which compares the elastic relaxation time $\lambda$ to the impact time $R_{0}/V_{0}$ (see § 2.1). While $De$ compares $\lambda$ to the inertio‑capillary time, $Wi$ is the more natural process‑time measure here and increases with the degree to which elastic stresses persist during impact WiandDe_Poole. Thus $Wi=0$ ($De=0$) recovers a Newtonian liquid, whereas $Wi\to\infty$ ($De\to\infty$) yields an elastic‑memory limit.
@@
The modeling choices in this work were made to isolate the physics of the liquid–to–elastic transition and enable direct comparison to the classical limits. The viscoelastic sphere obeys Oldroyd‑B, which neglects finite microstructure extensibility and shear‑thinning; this is acceptable for our flow history but can be systematically relaxed with constitutive models that incorporate finite extensibility and rate‑dependent viscosity Snoeijer_etal_2020, Mckinley_exten_flow, Thien_1977, tanner2000. The substrate is non‑contacting, so the thin gas layer is present but not explicitly resolved with lubrication and wetting dynamics; prior work shows that air cushioning, skating on a gas film, and nanoscale first contact depend sensitively on slip, compressibility and rarefaction smith-2003-jfm, thoroddsen-2003-jfm, thoroddsen_etoh_takehara_ootsuka_hatsuki_2005, kolinski-2012-prl, vdveen-2012-pre, Driscoll2011, li-2015-jfm, chubynsky-2020-prl, garcia2024skating, langley2017impact, zhang-2021-prf, sprittles2024gas. We stress that in the present work, the substrate is assumed to remain perfectly rigid. Small elastic deformations of the substrate can, however, modify the intervening gas layer Langley2020, which in turn may alter the impact dynamics and the measured force signature of the impacting sphere.Ma Furthermore, the onset of contact will modify the shear stress at the interface - including effects such as adhesion - that can modify the peak force, and will alter the stresses upon rebound. Finally, a small background viscosity renders the ($De\to\infty$) limit to be of Kelvin–Voigt type rather than ideally elastic; classical analyses quantify how viscoelastic dissipation and elastic waves perturb Hertzian impact pao1955extension, hunterEnergyAbsorbedElastic1957. These assumptions do not alter the governing exponents (Wagner versus Hertz) but they may affect quantitative prefactors and very‑short‑time, near‑contact details.
Several direct extensions can sharpen and generalize these results. Experiments with soft hydrogel/elastomer beads at moderate $V_{0}$ can probe the crossover regime and test the full $F(t)$ waveform (peak magnitude, rise time, symmetry/skewness), leveraging recent studies on soft‑solid and gel impacts and elastohydrodynamic bouncing mitra2021bouncing, jalaal-2018-jfm, luu-2009-jfm, martouzet2021dynamic, chevy2012liquid, aguero2022impact, bertinSimilaritySolutionsElastohydrodynamic2024, and using established force‑measurement protocols from liquid‑drop impacts Mitchell2019, Li2014, Zhang2017, zhangImpactForcesWater2022, sanjayRoleViscosityDrop2025. Furthermore, incorporating a thin‑gas lubrication model with dynamic wetting (including slip and, if needed, rarefaction/compressibility) will resolve when and how contact initiates and how this feeds back on the very‑early‑time force zheng2021air, smith-2003-jfm, kolinski-2012-prl, vdveen-2012-pre, li-2015-jfm, Driscoll2011, Mandre_etalPRL_2019, chubynsky-2020-prl, garcia2024skating, sprittles2024gas. Linking impact memory with lubricated‑impact and wetting transitions will place the near‑contact force history on firmer ground snoeijer-2013-arfm, Richard2000, sharma-2021-jfm, sanjay_chantelot_lohse_2023, chantelot_lohse_2021.
@@
All codes used in this work is available as an open-source repository at https://github.com/comphy-lab/Soft-Sphere-Impacts.

### v3

@@ 4.3 Predictive Interpolating Model for Maximum Impact Force
For intermediate conditions ($El\sim\mathcal{O}\left(1\right)$), the sphere’s deformation and the fluid’s inertia both contribute, and the peak force deviates from either pure Hertz or Wagner scaling alone. We therefore express the dimensionless peak force as a weighted transition between the two asymptotic contributions following the approach of Sanjay & Lohse (2025)1,
@@
with a smooth transition function $\text{f}\left(El\right)$ based on tanh function defined as
$\text{f}(El)=\frac{1+\text{tanh}(\frac{El-a}{b})}{2},$ (32)
where the $We$-dependant parameters $a\approx 1.3$ for $We\to 1$ and $\approx 0.5$ for $We\gtrsim 10$ and $b\approx 1.4$ for $We\to 1$ and $\approx 0.2$ for $We\gtrsim 10$ (see 25 for details of this fit). Here, $a$ measures the critical $El$ at transition from Wagner’s to Hertz’s scaling (fig.4). The width of this transition is indicated by $b$. The remaining coefficients in eq. (31) are fixed a priori: the Hertz prefactor 5.3 follows directly from the elastic analysis above (without fitting), while the constants 3.24 (Wagner plateau) and 3.2 (the $We^{-1}$ capillary correction) are taken from the Newtonian impact model of Sanjay & Lohse (2025)1. This construction ensures a continuous interpolation between the Hertz and Wagner limits. In spirit it follows the additive scaling approach of Sanjay & Lohse (2025)1 for drop impacts, but unlike their model – which includes separate viscous regimes – here only the two primary regimes (elastic vs. inertial) are needed. The resulting formula smoothly bridges the two asymptotes and correctly reproduces the peak-force scaling in both limits (this predictive curve is plotted in fig. 4a for comparison).
@@ 5 Influence of Elastic Stress Relaxation
The results in § 3 established the two asymptotic force scalings for impacts with permanent memory ($De\to\infty$): a Wagner plateau at small elasticity number and a Hertz law at large elasticity number (figure 4). These two limits also bound impacts when the material memory is finite: relaxing the memory shifts the response continuously from solid‑like to liquid‑like, with the transition controlled by the non‑dimensional relaxation time. We quantify the memory effect with the Weissenberg number, $Wi$ (eq. 8), which compares the elastic relaxation time $\lambda$ to the impact time $R_{0}/V_{0}$ (see § 2.1). While $De$ compares $\lambda$ to the inertio‑capillary time, $Wi$ is the more natural process‑time measure here and increases with the degree to which elastic stresses persist during impact 49. Thus $Wi=0$ ($De=0$) recovers a Newtonian liquid, whereas $Wi\to\infty$ ($De\to\infty$) yields an elastic‑memory limit.
@@
The modeling choices in this work were made to isolate the physics of the liquid–to–elastic transition and enable direct comparison to the classical limits. The viscoelastic sphere obeys Oldroyd‑B, which neglects finite microstructure extensibility and shear‑thinning; this is acceptable for our flow history but can be systematically relaxed with constitutive models that incorporate finite extensibility and rate‑dependent viscosity 63, 40, 67, 66. The substrate is non‑contacting, so the thin gas layer is present but not explicitly resolved with lubrication and wetting dynamics; prior work shows that air cushioning, skating on a gas film, and nanoscale first contact depend sensitively on slip, compressibility and rarefaction 61, 70, 69, 29, 73, 19, 33, 14, 21, 31, 77, 64. We stress that in the present work, the substrate is assumed to remain perfectly rigid. Small elastic deformations of the substrate can, however, modify the intervening gas layer 32, which in turn may alter the impact dynamics and the measured force signature of the impacting sphere.Ma Furthermore, the onset of contact will modify the shear stress at the interface - including effects such as adhesion - that can modify the peak force, and will alter the stresses upon rebound. Finally, a small background viscosity renders the ($De\to\infty$) limit to be of Kelvin–Voigt type rather than ideally elastic; classical analyses quantify how viscoelastic dissipation and elastic waves perturb Hertzian impact 47, 23. These assumptions do not alter the governing exponents (Wagner versus Hertz) but they may affect quantitative prefactors and very‑short‑time, near‑contact details.
Several direct extensions can sharpen and generalize these results. Experiments with soft hydrogel/elastomer beads at moderate $V_{0}$ can probe the crossover regime and test the full $F(t)$ waveform (peak magnitude, rise time, symmetry/skewness), leveraging recent studies on soft‑solid and gel impacts and elastohydrodynamic bouncing 42, 24, 37, 39, 13, 2, 30, and using established force‑measurement protocols from liquid‑drop impacts 41, 34, 75, 76, 56. Furthermore, incorporating a thin‑gas lubrication model with dynamic wetting (including slip and, if needed, rarefaction/compressibility) will resolve when and how contact initiates and how this feeds back on the very‑early‑time force 78, 61, 29, 73, 33, 19, 38, 14, 21, 64. Linking impact memory with lubricated‑impact and wetting transitions will place the near‑contact force history on firmer ground 62, 54, 58, 55, 10.
@@
All codes used in this work is available as an open-source repository at https://github.com/comphy-lab/Soft-Sphere-Impacts 25.
@@ Acknowledgements
Notes and references
  • V. Sanjay and D. Lohse (2025) Unifying theory of scaling in drop impact: forces and maximum spreading diameter. Phys. Rev. Lett. 134, pp. 104003. External Links: Document Cited by: §1, §1, §4.2, §4.2, §4.3, §4.3.
  • E. A. Agüero, L. F. L. Alventosa, D. M. Harris, and C. A. Galeano-Rios (2022) Impact of a rigid sphere onto an elastic membrane. Proc. R. Soc. Lond. A 478 (2266), pp. 20220340. Cited by: Conclusions & outlook.
  • M. Ahmad, M. Schatz, and M. Casey (2013) Experimental investigation of droplet size influence on low pressure steam turbine blade erosion. Wear 303, pp. 83–86. External Links: Document Cited by: §1.
  • B. Andreotti, Y. Forterre, and O. Pouliquen (2013) Granular media: between fluid and solid. Cambridge University Press. Cited by: §1.
  • V. B. Bergeron, D. B. Bonn, J. M. Martin, and V. L. Louis (2000) Controlling droplet deposition with polymer additives. Nature 405, pp. 772–775. External Links: Document Cited by: §1.
  • R.B. Bird, P.J. Dotson, and N.L. Johnson (1980) Polymer solution rheology based on a finitely extensible bead—spring chain model. J. Non-Newtonian Fluid Mech. 7 (2), pp. 213–235. External Links: ISSN 0377-0257, Document Cited by: §2.1.
  • J. U. Brackbill, D. B. Kothe, and C. Zemach (1992) A continuum method for modeling surface tension. J. Comput. Phys. 100 (2), pp. 335–354. Cited by: §2.2.
  • H. Hertz (1882) Ueber die berührung fester elastischer körper.. Journal für die reine und angewandte Mathematik 1882, pp. 156–171 (en). External Links: Link, Document, ISSN 0075-4102, 1435-5345 Cited by: §1, §4.1.
  • P. Calvert (2001) Inkjet printing for materials and devices. Chem. Mater. 13, pp. 3299–3305. External Links: Document Cited by: §1.
  • P. Chantelot and D. Lohse (2021) Drop impact on superheated surfaces: short-time dynamics and transition to contact. J. Fluid Mech. 928, pp. A36. Cited by: Conclusions & outlook.
  • D. Chen, Q. Wen, P. Janmey, J. Crocker, and A. Yodh (2010) Rheology of soft materials. Annu. Rev. Condens. Matter Phys 1, pp. 301–22. External Links: Document Cited by: §1.
  • X. Cheng, T.-P. Sun, and L. Gordillo (2022) Drop impact dynamics: impact force and stress distributions. Annu. Rev. Fluid Mech. 54, pp. 57–81. Cited by: §1, §1, §4.2.
  • F. Chevy, A. Chepelianskii, D. Quéré, and E. Raphaël (2012) Liquid Hertz contact: softness of weakly deformed drops on non-wetting substrates. Europhys. Lett. 100 (5), pp. 54002. Cited by: Conclusions & outlook.
  • M. V. Chubynsky, K. I. Belousov, D. A. Lockerby, and J. E. Sprittles (2020) Bouncing off the walls: the influence of gas-kinetic and van der Waals effects in drop impact. Phys. Rev. Lett. 124, pp. 084501. Cited by: Conclusions & outlook, Conclusions & outlook.
  • C. Clasen, J. Eggers, M. A. Fontelos, J. Li, and G. H. Mckinley (2006) The beads-on-string structure of viscoelastic threads. J. Fluid Mech. 556, pp. 283–308. External Links: Document Cited by: §1.
  • R. Cross (1999) The bounce of a ball. Am. J. Phys. 67 (3), pp. 222–227. External Links: Document, ISSN 0002-9505 Cited by: §1.
  • B. Derby (2012) Printing and prototyping of tissues and scaffolds. Science 338, pp. 921–6. External Links: Document Cited by: §1.
  • A. K. Dixit, A. Oratis, K. Zinelis, D. Lohse, and V. Sanjay (2025) Viscoelastic Worthington jets and droplets produced by bursting bubbles. J. Fluid Mech. 1010, pp. A2. External Links: ISSN 0022-1120, 1469-7645, Document Cited by: §2.1, §2.2.
  • M. M. Driscoll and S. R. Nagel (2011) Ultrafast interference imaging of air in splashing dynamics. Phys. Rev. Lett. 107 (15), pp. 154502. Cited by: Conclusions & outlook, Conclusions & outlook.
  • J. Eggers, M. A. Fontelos, C. Josserand, and S. Zaleski (2010) Drop dynamics after impact on a solid wall: theory and simulations. Phys. Fluids 22 (6), pp. 062101. Cited by: §4.2.
  • P. García-Geijo, G. Riboux, and J. Gordillo (2024) The skating of drops impacting over gas or vapour layers. J. Fluid Mech. 980, pp. A35. Cited by: Conclusions & outlook, Conclusions & outlook.
  • L. He, L. Ding, B. Li, W. Mu, P. Li, and F. Liu (2021) Optimization strategy to inhibit droplets rebound on Pathogen-Modified hydrophobic surfaces. ACS Appl. Mater. Interfaces 13 (32), pp. 38018–38028. Cited by: §1.
  • S. C. Hunter (1957) Energy absorbed by elastic waves during impact. J. Mech. Phys. Solids 5 (3), pp. 162–171. External Links: ISSN 0022-5096, Document Cited by: §1, Conclusions & outlook.
  • M. Jalaal, C. Seyfert, B. Stoeber, and N. J. Balmforth (2018) Gel-controlled droplet spreading. J. Fluid Mech. 837, pp. 115–128. Cited by: Conclusions & outlook.
  • S. Jana and V. Sanjay (2026) Code repository: Soft Sphere Impacts. Note: https://github.com/comphy-lab/Soft-Sphere-Impacts Cited by: §4.3, Data availability.
  • C. Josserand and S. T. Thoroddsen (2016) Drop impact on a solid surface. Annu. Rev. Fluid Mech. 48, pp. 365–391. Cited by: §4.2.
  • J. Kim (2007) Spray cooling heat transfer: the state of the art. Int. J. Heat Fluid Flow 28 (4), pp. 753–767. External Links: Document, ISSN 0142-727X Cited by: §1.
  • J. M. Kolinski, L. Mahadevan, and S. M. Rubinstein (2014) Drops can bounce from perfectly hydrophilic surfaces. Europhys. Lett. 108, pp. 24001. Cited by: §2.2.
  • J. M. Kolinski, S. M. Rubinstein, S. Mandre, M. P. Brenner, D. A. Weitz, and L. Mahadevan (2012) Skating on a film of air: drops impacting on a surface. Phys. Rev. Lett. 108, pp. 074503. Cited by: Conclusions & outlook, Conclusions & outlook.
  • V. Bertin (2024) Similarity solutions in elastohydrodynamic bouncing. J. Fluid Mech. 986, pp. A13. External Links: Link, Document Cited by: §1, §1, Conclusions & outlook.
  • K. Langley, E. Q. Li, and S. T. Thoroddsen (2017) Impact of ultra-viscous drops: air-film gliding and extreme wetting. J. Fluid Mech. 813, pp. 647–666. Cited by: Conclusions & outlook.
  • K. R. Langley, A. A. Castrejón-Pita, and S. T. Thoroddsen (2020) Droplet impacts onto soft solids entrap more air. Soft Matter 16, pp. 5702–5710. External Links: Document Cited by: Conclusions & outlook.
  • E. Q. Li, I. U. Vakarelski, and S. T. Thoroddsen (2015) Probing the nanoscale: the first contact of an impacting drop. J. Fluid Mech. 785, pp. R2. Cited by: Conclusions & outlook, Conclusions & outlook.
  • J. Li, B. Zhang, P. Guo, and Q. Lv (2014) Impact force of a low speed water droplet colliding on a solid surface. J. Appl. Phys. 116 (21), pp. 214903. Cited by: Conclusions & outlook.
  • J. Li and D. Mooney (2016) Designing hydrogels for controlled drug delivery. Nature Rev. Mater. 1, pp. 16071. External Links: Document Cited by: §1.
  • D. Lohse (2022) Fundamental fluid dynamics challenges in inkjet printing. Annu. Rev. Fluid Mech. 54, pp. 349–382. Cited by: §1.
  • L.-H. Luu and Y. Forterre (2009) Drop impact of yield-stress fluids. J. Fluid Mech. 632, pp. 301–327. Cited by: Conclusions & outlook.
  • S. Mandre, M. Mani, and M. P. Brenner (2009) Precursors to splashing of liquid droplets on a solid surface. Phys. Rev. Lett. 102, pp. 134502. External Links: Document, Link Cited by: §4.2, Conclusions & outlook.
  • G. Martouzet, L. Jørgensen, Y. Pelet, A.-L. Biance, and C. Barentin (2021) Dynamic arrest during the spreading of a yield stress fluid drop. Phys. Rev. Fluids 6 (4), pp. 044006. Cited by: Conclusions & outlook.
  • G. H. McKinley and T. Sridhar (2002) Filament-stretching rheometry of complex fluids. Annu. Rev. Fluid Mech. 34, pp. 375–415. External Links: Document, ISSN 1545-4479 Cited by: §2.1, Conclusions & outlook.
  • B. R. Mitchell, J. C. Klewicki, Y. P. Korkolis, and B. L. Kinsey (2019) The transient force profile of low-speed droplet impact: measurements and model. J. Fluid Mech. 867, pp. 300–322. Cited by: Conclusions & outlook.
  • S. Mitra, Q. Vo, and T. Tran (2021) Bouncing-to-wetting transition of water droplets impacting soft solids. Sodt Matter. Cited by: Conclusions & outlook.
  • M. A. Nearing, J. M. Bradford, and R. D. Holtz (1986) Measurement of force vs. time relations for waterdrop impact. Soil Sci. Soc. Am. J. 50, pp. 1532–1536. Cited by: §1.
  • M. J. Negus, M. R. Moore, J. M. Oliver, and R. Cimpeanu (2021) Droplet impact onto a spring-supported plate: analysis and simulations. J. Eng. Math. 128 (1), pp. 3. Cited by: §4.2.
  • M. Negus (2022) Modelling droplet impact in fluid-structure interaction problems. Ph.D. Thesis, University of Oxford. Cited by: §4.2.
  • J. G. Oldroyd and A. H. Wilson (1950) On the formulation of rheological equations of state. Proc. R. Soc. Lond. A 200 (1063), pp. 523–541. External Links: Document Cited by: §2.1.
  • Y.-H. Pao (1955) Extension of the Hertz theory of impact to the viscoelastic case. J. Appl. Phys. 26 (9), pp. 1083–1088. Cited by: §1, Conclusions & outlook.
  • J. Philippi, P.-Y. Lagrée, and A. Antkowiak (2016) Drop impact on a solid surface: short-time self-similarity. J. Fluid Mech. 795, pp. 96–135. Cited by: §1.
  • R. J. Poole (2012) The Deborah and Weissenberg numbers. Rheol. Bull. 53 (2), pp. 32–39. External Links: Link Cited by: §5.
  • S. Popinet and collaborators (2013) Basilisk C: volume of fluid method. Note: http://basilisk.fr (Last accessed: August 23, 2023) Cited by: §2.2.
  • S. Popinet (2018) Numerical models of surface tension. Annu. Rev. Fluid Mech. 50, pp. 49–75. Cited by: §2.2.
  • S. Popinet (2009) An accurate adaptive solver for surface-tension-driven interfacial flows. J. Comput. Phys. 228 (16), pp. 5838–5866. External Links: Document, ISSN 0021-9991, Link Cited by: §2.2, §2.2.
  • S. Popinet (2015) A quadtree-adaptive multigrid solver for the serre–green–naghdi equations. J. Comput. Phys. 302, pp. 336–358. External Links: Document, ISSN 0021-9991 Cited by: §2.2.
  • D. Richard and D. Quéré (2000) Bouncing water drops. Europhys. Lett. 50 (6), pp. 769–775. Cited by: Conclusions & outlook.
  • V. Sanjay, P. Chantelot, and D. Lohse (2023) When does an impacting drop stop bouncing?. J. Fluid Mech. 958, pp. A26. External Links: Document Cited by: Conclusions & outlook.
  • V. Sanjay, B. Zhang, C. Lv, and D. Lohse (2025) The role of viscosity on drop impact forces on non-wetting surfaces. J. Fluid Mech. 1004, pp. A6. External Links: Document Cited by: §1, §1, Fig. 2, Fig. 2, §3, §3, §4.2, §4.2, Conclusions & outlook.
  • V. Sanjay (2022) Viscous free-surface flows. PhD Thesis - Research UT, graduation UT, University of Twente, University of Twente, Netherlands. External Links: Document Cited by: §2.2, §3.
  • P. K. Sharma and H. N. Dixit (2021) Regimes of wettability-dependent and wettability-independent bouncing of a drop on a solid surface. J. Fluid Mech. 908, pp. A37. Cited by: Conclusions & outlook.
  • S. Shiri and J. C. Bird (2017) Heat exchange between a bouncing drop and a superhydrophobic substrate. Proc. Natl. Acad. Sci. USA 114 (27), pp. 6930–6935. External Links: Document Cited by: §1.
  • F. R. Smith, C. Nicloux, and D. Brutin (2018) Influence of the impact energy on the pattern of blood drip stains. Phys. Rev. Fluids 3, pp. 013601. External Links: Document, Link Cited by: §1.
  • F. T. Smith, L. Li, and G. X. Wu (2003) Air cushioning with a lubrication/inviscid balance. J. Fluid Mech. 482, pp. 291–318. Cited by: Conclusions & outlook, Conclusions & outlook.
  • J. H. Snoeijer and B. Andreotti (2013) Moving contact lines: scales, regimes, and dynamical transitions. Annu. Rev. Fluid Mech. 45, pp. 269–292. Cited by: Conclusions & outlook.
  • J. H. Snoeijer, A. Pandey, M. A. Herrada, and J. Eggers (2020) The relationship between viscoelasticity and elasticity. Proc. R. Soc. A 476 (2243), pp. 20200419. External Links: Document Cited by: §1, §2.1, §2.1, §2.1, Conclusions & outlook.
  • J. E. Sprittles (2024) Gas microfilms in droplet dynamics: when do drops bounce?. Annu. Rev. Fluid Mech. 56, pp. 91–118. Cited by: Conclusions & outlook, Conclusions & outlook.
  • S. Tamim and J. Bostwick (2021) Oscillations of a soft viscoelastic drop. npj Microgravity 7, pp. 42. External Links: Document Cited by: §1.
  • R.I. Tanner (2000) Engineering rheology. Oxford Engineering Science Series, OUP Oxford. External Links: ISBN 9780191590160, Link Cited by: §2.1, Conclusions & outlook.
  • N. P. Thien and R. I. Tanner (1977) A new constitutive equation derived from network theory. J. Non-Newton. Fluid 2 (4), pp. 353–365. External Links: Document, ISSN 0377-0257 Cited by: §2.1, Conclusions & outlook.
  • J. Thompson and J. Rothstein (2007) The atomization of viscoelastic fluids in flat-fan and hollow-cone spray nozzles. J. Non-Newtonian Fluid Mech. 147, pp. 11–22. External Links: Document Cited by: §1.
  • S. T. Thoroddsen, T. G. Etoh, K. Takehara, N. Ootsuka, and Y. Hatsuki (2005) The air bubble entrapped under a drop impacting on a solid surface. J. Fluid Mech. 545, pp. 203–212. External Links: Document Cited by: Conclusions & outlook.
  • S. T. Thoroddsen, T. G. Etoh, and K. Takehara (2003) Air entrapment under an impacting drop. J. Fluid Mech. 478, pp. 125–134. Cited by: Conclusions & outlook.
  • Y. Tirupataiah and G. Sundararajan (1991) A dynamic indentation technique for the characterization of the high strain rate plastic flow behaviour of ductile metals and alloys. J. Mech. Phys. Solids 39 (2), pp. 243–271. External Links: Document, ISSN 0022-5096 Cited by: §1.
  • G. Tryggvason, R. Scardovelli, and S. Zaleski (2011) Direct Numerical Simulations of Gas–Liquid Multiphase Flows. Cambridge University Press. Cited by: §2.2.
  • R. C. A. van der Veen, T. Tran, D. Lohse, and C. Sun (2012) Direct measurements of air layer profiles under impacting droplets using high-speed color interferometry. Phys. Rev. E 85, pp. 026315. Cited by: Conclusions & outlook, Conclusions & outlook.
  • H. Wagner (1932) Über Stoß- und Gleitvorgänge an der Oberfläche von Flüssigkeiten. Z. Angew. Math. Mech. 12 (4), pp. 193–215. Cited by: §1.
  • B. Zhang, J. Li, P. Guo, and Q. Lv (2017) Experimental studies on the effect of reynolds and weber numbers on the impact forces of low-speed droplets colliding with a solid surface. Exp. Fluids 58 (9), pp. 1–12. Cited by: Conclusions & outlook.
  • B. Zhang, V. Sanjay, S. Shi, Y. Zhao, C. Lv, and D. Lohse (2022) Impact forces of water drops falling on superhydrophobic surfaces. Phys. Rev. Lett. 129, pp. 104501. Cited by: §1, §1, Fig. 2, Fig. 2, §3, §3, §3, §4.2, §4.2, Conclusions & outlook.
  • L. Zhang, T. Soori, A. Rokoni, A. Kaminski, and Y. Sun (2021) Thin film instability driven dimple mode of air film failure during drop impact on smooth surfaces. Phys. Rev. Fluids 6, pp. 044002. Cited by: Conclusions & outlook.
  • S. Zheng, S. Dillavou, and J. M. Kolinski (2021) Air mediates the impact of a compliant hemisphere on a rigid smooth surface. Soft Matter 17 (14), pp. 3813–3819. Cited by: §2.2, Conclusions & outlook.